# Unusual critical points between atomic insulating phases

### v1

@@ Appendix C Velocity Anisotropy in QED3
In this section, we will analyze the fate of velocity anisotropy in QED3 under the renormalization group, in the large $N_{f}$ limit, similar to the analysis done in Franz et al. (2002); Vafek et al. (2002); Hermele et al. (2005, 2007). To begin, we have the Euclidean Lagrangian
@@ C.3 Renormalization
The above suggests that the QED3 fixed point is not destabilized by velocity anisotropies, except for the two marginal anisotropic perturbations above. However, we will demonstrate that the two marginal operators $\Gamma_{2}$ and $\Gamma_{4}$ are redundant operators, and therefore, do not affect the conformal fixed point. In more detail, a redundant operator is any operator in the action that can be removed by a continuous definition of the fields Weinberg (2005),
@@
Thus, a redundant operator is equivalently one that vanishes on the equations of motion, $\frac{\delta\mathcal{L}}{\delta\Psi}=0$. Strictly speaking, the renormalization group behavior of a redundant operator is scheme dependent and not universal Wegner (1974). In our renormalization group flow, both $\Gamma_{2}$ and $\Gamma_{4}$ are marginal as they arise from a constant rescaling (or reparametrization) of $\Psi$ at the fixed point. Specifically, $\Gamma_{2}$ can be eliminated from the action at the fixed point theory by redefining $\Psi\rightarrow{\Psi}-\frac{\Gamma_{2}}{2}\mu{\Psi}$, followed by a field renormalization, while $\Gamma_{4}$ can be eliminated by redefining ${\Psi}\rightarrow\Psi+i\frac{\Gamma_{4}}{2}\mu\gamma^{0}\Psi$ and then rescaling the time coordinate.
@@ Appendix D LSM Anomaly Matching
We follow the approach outlined in Ye et al. (2022). Recall that from Eq. (43) in the main text we must match the anomalies as elements of the group cohomology $H^{4}(G_{UV},U(1)_{T})$. The $U(1)$ coefficients complicate the process, but for the groups we are considering, $H^{4}(G_{UV},U(1)_{T})$ is just a product of $\mathbb{Z}_{2}$ factors. Then, it is much simpler to calculate the IR and UV anomalies as elements of $H^{4}(G_{UV},\mathbb{Z}_{2})$. This equates to viewing
@@
In particular $\exp(i\pi\bullet)$ maps the $\mathbb{Z}_{2}$ valued class $\Omega[G]$ into the $U(1)$ valued anomaly $\eta[G]$. More details can be found in Ye et al. (2022).
@@ D.1 The form of $\eta$ and $\Omega$ for $G_{UV/IR}$
Now in the IR, we have $N_{f}=4$ QED3 with $G_{IR}={SO(6)\times U(1)_{top}}/{\mathbb{Z}_{2}}$. However, there is not yet a known form of the anomaly $G_{IR}$ in terms of characteristic classes for the $G_{IR}$ anomaly for $N_{f}=4$ QED3. Instead, we will consider a slightly simplified description of the anomaly by coupling $SO(6)$ and $U(1)_{top}=SO(2)$ gauge fields, considering the QED3 CFT as a Stiefel liquid Zou et al. (2021), which provides a dual description of the theory in terms of a nonlinear sigma model. Allowing time reversal, we must then consider an $O(6)_{T}\times O(2)_{T}$ gauge bundle, where the subscript $T$ indicates that improper rotations in $O(2)$ or $O(6)$ must be composed with a spacetime orientation reversal symmetry. Mathematically, this constrains
@@
in terms of the Stiefel-Whitney classes of the tangent, $O(6)$, and $O(2)$ bundles. The ’t Hooft anomaly is then characterized by an SPT bulk, Zou et al. (2021)

### v2

@@ Appendix C Velocity Anisotropy in QED3
In this section, we will analyze the fate of velocity anisotropy in QED3 under the renormalization group, in the large $N_{f}$ limit, similar to the analysis done in [24, 63, 32, 33]. To begin, we have the Euclidean Lagrangian
@@ C.3 Renormalization
The above suggests that the QED3 fixed point is not destabilized by velocity anisotropies, except for the two marginal anisotropic perturbations above. However, we will demonstrate that the two marginal operators $\Gamma_{2}$ and $\Gamma_{4}$ are redundant operators, and therefore, do not affect the conformal fixed point. In more detail, a redundant operator is any operator in the action that can be removed by a continuous definition of the fields [69],
@@
Thus, a redundant operator is equivalently one that vanishes on the equations of motion, $\frac{\delta\mathcal{L}}{\delta\Psi}=0$. Strictly speaking, the renormalization group behavior of a redundant operator is scheme dependent and not universal [68]. In our renormalization group flow, both $\Gamma_{2}$ and $\Gamma_{4}$ are marginal as they arise from a constant rescaling (or reparametrization) of $\Psi$ at the fixed point. Specifically, $\Gamma_{2}$ can be eliminated from the action at the fixed point theory by redefining $\Psi\rightarrow{\Psi}-\frac{\Gamma_{2}}{2}\mu{\Psi}$, followed by a field renormalization, while $\Gamma_{4}$ can be eliminated by redefining ${\Psi}\rightarrow\Psi+i\frac{\Gamma_{4}}{2}\mu\gamma^{0}\Psi$ and then rescaling the time coordinate.
@@ Appendix D LSM Anomaly Matching
We follow the approach outlined in [72]. Recall that from Eq. (43) in the main text we must match the anomalies as elements of the group cohomology $H^{4}(G_{UV},U(1)_{T})$. The $U(1)$ coefficients complicate the process, but for the groups we are considering, $H^{4}(G_{UV},U(1)_{T})$ is just a product of $\mathbb{Z}_{2}$ factors. Then, it is much simpler to calculate the IR and UV anomalies as elements of $H^{4}(G_{UV},\mathbb{Z}_{2})$. This equates to viewing
@@
In particular $\exp(i\pi\bullet)$ maps the $\mathbb{Z}_{2}$ valued class $\Omega[G]$ into the $U(1)$ valued anomaly $\eta[G]$. More details can be found in [72].
@@ D.1 The form of $\eta$ and $\Omega$ for $G_{UV/IR}$
Now in the IR, we have $N_{f}=4$ QED3 with $G_{IR}={SO(6)\times U(1)_{top}}/{\mathbb{Z}_{2}}$. However, there is not yet a known form of the anomaly $G_{IR}$ in terms of characteristic classes for the $G_{IR}$ anomaly for $N_{f}=4$ QED3. Instead, we will consider a slightly simplified description of the anomaly by coupling $SO(6)$ and $U(1)_{top}=SO(2)$ gauge fields, considering the QED3 CFT as a Stiefel liquid [76], which provides a dual description of the theory in terms of a nonlinear sigma model. Allowing time reversal, we must then consider an $O(6)_{T}\times O(2)_{T}$ gauge bundle, where the subscript $T$ indicates that improper rotations in $O(2)$ or $O(6)$ must be composed with a spacetime orientation reversal symmetry. Mathematically, this constrains
@@
in terms of the Stiefel-Whitney classes of the tangent, $O(6)$, and $O(2)$ bundles. The ’t Hooft anomaly is then characterized by an SPT bulk, [76]